OpenStudy (anonymous):

Please Help! I need an explanation for how to do this... Step-by-step directions would be great! Simplify. Assume the divisor is not zero. \[(8x^2y^2 – 4xy^2) \over 4xy\]

OpenStudy (anonymous):

Answer choices include: A. xy – 1 B. 2xy C. 2xy – 1 D. 2xy – y

OpenStudy (asnaseer):

First look at the numerator of this fraction - can you see a common constant number that will divide into both terms in the numerator?

OpenStudy (anonymous):

I believe four or two....

OpenStudy (asnaseer):

4 is indeed the correct value, so now you can write your fraction as:\[\frac{8x^2y^2 – 4xy^2}{4xy}=\frac{4(2x^2y^2 – xy^2)}{4xy}\]Now look at the \(x\) terms in the numerator - what power of \(x\) is common to both terms in the numerator?

OpenStudy (anonymous):

2 or 1... I am not quite sure. :(

OpenStudy (anonymous):

I believe the final answer is D. 2xy - y. Is it correct? @asnaseer

OpenStudy (asnaseer):

yes :)

OpenStudy (anonymous):

Really? Wow! Thank you! I have two more problems, may you help with those as well?

OpenStudy (asnaseer):

so we can write this as:\[\frac{8x^2y^2 – 4xy^2}{4xy}=\frac{4(2x^2y^2 – xy^2)}{4xy}=\frac{4x(2xy^2 – y^2)}{4xy}=\frac{4xy(2xy – y)}{4xy}=\frac{\cancel{4xy}(2xy – y)}{\cancel{4xy}}\]\[=2xy-y\]

OpenStudy (asnaseer):

yw :) Please post each question separately and I am sure there will be plenty of people that will try and help you out (if I am busy with other things).

OpenStudy (anonymous):

Yes, I have a similar equation written in my notebook, but your equation definitely looks nicer. :)

OpenStudy (anonymous):

And yes, I have posted the next question. May I tag you in the question?

OpenStudy (anonymous):

@asnaseer ^

OpenStudy (asnaseer):

sure - but, as I said, I am not always free but will try my best :)

OpenStudy (anonymous):

Well, if you cannot help, then do not worry. :)